OpenStudy (anonymous):

PLEASE HELP!you draw 2 cards from a well-shuffled deck of decks cards. after you draw the first card you do not replace it in the deck. what is the probability of p (get ace of spades on first draw and get queen of hearts on second draw)?

OpenStudy (anonymous):

There's only one ace of spades and one queen of hearts, so the probability of drawing as ace of spades is \(\dfrac{1}{52}\). Since you don't replace the first card drawn, there are 51 cards left, one of which is the desired queen of hearts; so, the probability of getting the queen of hearts on the second draw is \(\dfrac{1}{51}\). The probability of drawing an ace of spades first and a queen of hearts second is thus \(\dfrac{1}{52}\cdot\dfrac{1}{51}\).

OpenStudy (anonymous):

@SithsAndGiggles would 1/2652 be correct?

OpenStudy (anonymous):

Yes